OpenStudy (anonymous):

Simplify the radical. 4 square root of 20

OpenStudy (anonymous):

would it be 400..

OpenStudy (zzr0ck3r):

\(4\sqrt{20}=4\sqrt{4*5}=4\sqrt{4}\sqrt{5}\)

OpenStudy (anonymous):

It cant be, these are the option, 5 to the square root of 16 5 to the square root of 8 8 to the square root of 5 16 to the square root of 5

OpenStudy (anonymous):

for \[4\sqrt{20}\] i got 17.88

OpenStudy (anonymous):

But that is not one of the options. @emily_wilson

OpenStudy (anonymous):

I see that

OpenStudy (anonymous):

Then what would it be?

OpenStudy (anonymous):